Write the slope-intercept equation for the line with the given slope that contains the given point.
m = 3/2; (7, 4)

Answer by stanbon(75887)

Write the slope-intercept equation for the line with the given slope that contains the given point.
m = 3/2; (7, 4)
-----
Form: y = mx + b
Substitute and solve for "b":
4 = (3/2)7 + b
b = (8/2)-(21/2) = -13/2
-----
Equation:
y = (3/2)x - (13/2)
=======================
Cheers,
Stan H.